OpenStudy (anonymous):

Which sets of data show a wave with the shortest wavelength? speed = 100 million m/s and frequency = 50 million Hz speed = 150 million m/s and frequency = 1500 Hz speed = 300 million m/s and frequency = 100 Hz

OpenStudy (anonymous):

I think it's c but im not a hundred percent sure

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your Welcome :) You Fan though :P

OpenStudy (matt101):

You can find the answer using the basic wave equation: v=fλ, where v is the speed, f is the frequency, and λ is the wavelength. In this case we're trying to solve for wavelength, so just rearrange the equation: λ=v/f. For A we have (10^8)/(5*10^7) = 2 m For B we have (1.5*10^8)/(1.5*10^3) = 10^5 m For C we have (3*10^8)/(10^2)=3*10^6 m You can see from this simple calculation that A has the shortest wavelength, so that's your answer.